# Entanglement-breaking channels are a quantum memory resource

### v1

@@ Abstract
Crucially, the maxima $\Omega_{C}(\bm{a},d)$ must be found explicitly within classical theory, either by exact computation or by obtaining an upper bound, whereas for the quantum value of $p(\bm{a}|\mathcal{I},\mathcal{E},d)$ an explicit model $(\mathcal{I},\mathcal{E})$ must be found. If the classical memory bound has been violated even in the presence of an EB channel $\mathcal{E}$, as in Eq. 7, then the system still acts as a quantum memory resource for the task of generating temporal correlations.
To achieve this goal, we consider the one-tick sequences [8, 25] of length $L$, defined as $\bm{a}_{\text{ot}}^{L}:=0\dots 01$ and having $\operatorname{DC}(\bm{a}^{L}_{\text{ot}})=L$, under the scenario $d=L-1$. As the case of $L=2$ is trivial, we consider the smallest non-trivial cases $L\geq 3$. For the classical model, analytical upper bounds can be computed via semidefinite programming (SDP) [29]. The case $L=3$ and $4$ were already solved in [29], and here we computed explicitly the case $L=5$ with the same method. In all these cases, the upper bounds computed via SDP coincide, up to the numerical precision, with the values obtained via the one-way model [8], which is conjectured to be optimal for all $d=L-1$ scenarios [8, 25]. We are therefore sure to have obtained the exact bounds for the cases $L=3,4,$ and $5$. An explicit quantum model violating these classical bounds can be obtained with the following construction. Consider the deterministic classical model on $m$ states given by $[T_{0}]_{i,i+1}=1$ for $i=1,\dots,m-1$, $[T_{1}]_{m,1}=1$, zeros for all other entries, and an initial state $\pi=(1,0,\dots,0)$. The idea is to approximate this model by emulating $m=d+1$ nearly-orthogonal virtual states with the available $d$-dimensional quantum system.
With $\zeta=\mathrm{e}^{2\pi\mathrm{i}/(m-1)}$, we define $\mathcal{E}$ using
@@
then let $\sigma_{n}=\left|\psi_{n}\right\rangle\left\langle\psi_{n}\right|$ and $E_{n+1}=\tfrac{d-1}{d}\sigma_{n}$ for $n=1,\dots,m-1$, and finally setting $\rho_{0}=\sigma_{m}=E_{1}=\left|0\right\rangle\left\langle 0\right|$. The instrument is defined as $\mathcal{I}_{a}(\rho)=K_{a}\rho K^{\dagger}_{a}$ with $K_{0}=\mathrm{diag}(1,\cdots,1,0)$ and $K_{1}=\mathbbm{1}-K_{0}$. This quantum model corresponds to isolating the $(d-1)$-dimensional subspace orthogonal to $\left|0\right\rangle$, assembling a harmonic complex equiangular tight frame (ETF) [31] of size $m-1$ within this subspace, then using it for the virtual states in the transitions of $T_{0}$. ETFs are useful for this task as they provide a generalization of orthonormal bases, saturating the Welch bounds for minimum overlaps [32]. As seen in Fig. 3, this quantum ETF construction clearly violates the exact classical bounds for $L=4$ and $5$, conclusively showing that general EB channels cannot be applied to characterize classical memory effects.
@@
The quantum model presented above provides an explicit violation of the classical bound. This is, however, not necessarily the maximum one can achieve in the quantum case. To investigate the quantum maximum with EB channels, we have also performed numerical optimization via gradient descent methods, obtaining lower bounds for the maxima $\Omega_{Q}(\bm{a},d)$. Results are summarized in Table 1. The quantum models found numerically slightly outperform the previous ETF construction. See App. C for more details regarding the optimization and App. D for the explicit models found.
@@
Conclusion and discussion.— Entanglement breaking channels are universally recognized as classical memories [1, 2, 3, 9, 16, 4, 10, 17, 12, 11], and, indeed they are classical under many points of view. Quite surprisingly, however, we found that there exist scenarios where they still act as genuine quantum memories. More precisely, we answer in the negative the following question: is a qudit passing through an entanglement-breaking channel equivalent to a classical $d$-it?
@@
To address this question, we have investigated a task where memory is the central resource, namely, the generation of correlations in time. Indeed, it has been shown that any temporal correlation can be reproduced by a classical system if enough memory is available [5, 26, 27], such that differences between classical and quantum temporal correlations only emerge if memory is bounded [6, 8, 25]. As classical memory is understood in terms of the number of internal states $d$ of the system, a fair comparison between classical and quantum theory must involve a quantum system of the same size. Differences between quantum and classical memories can be found by investigating the most general set of memory operations allowed in each theory, then establishing bounds on their capabilities.
The violations of classical bounds we have found showcase the importance of investigating classical memory effects in terms of manifestly classical descriptions, instead of relying on quantum-inspired analogies, as otherwise unexpected quantum effects may persist. As we have shown, the common presumption that EB channels act as a classical resource carries additional assumptions, which could be rather strong; see discussion on App. B. Furthermore, our results emphasize that, in the context of temporal correlations, it is essential to look beyond qubits and two-time scenarios, as otherwise important distinctions between classical and quantum systems are not apparent, as also recognized in Ref. [17]; see also App. D.
@@
Our findings inspire several future research directions. Understanding the minimal requirements for an EB channel to truly act as a classical memory could offer insights into the origins of the quantum advantage in the context of temporal correlations, allowing a more definitive understanding of classical and quantum memory effects in non-Markovian processes; see also App. B. Furthermore, since the quantum advantage seen in this work stems from the non-orthogonality of the states $\sigma_{i}$, an investigation of the quantum upper bounds in the presence of EB channels could lead to new insights into the optimal encoding of classical information within quantum systems for various sequential tasks, with potential applications to, e.g., quantum random-access codes [33, 34, 35, 36] or quantum compression of classical predictive models for stochastic processes [37, 38, 39].
@@
Naturally, there are many distinct features one could wish from a “quantum memory”, e.g., preserving coherence or entanglement [1, 9, 10], or enabling non-classical multi-time statistics [7, 17]. If the quantum memory is required to preserve entanglement, then EB channels are obviously useless as a quantum resource. However, quantum effects are not restricted to spatial correlations, and in considering the temporal character inherent to any information processing task, we are required to also consider the fundamental differences between classical and quantum temporal correlations. Our results indicate a more careful and detailed investigation of quantum memory resources is warranted, such that their precise origin and advantage over classical memories can be properly understood, quantified, and successfully applied in emergent quantum technologies.
@@
Acknowledgements.— We thank Simon Milz and Philip Taranto for valuable discussions. This work is supported by the Austrian Science Fund (FWF) through projects ZK 3 (Zukunftskolleg) and F7113 (BeyondC). This work is supported by the Ministry of Science and Technology, Taiwan, (Grants No. MOST 111-2917-I-564-005, 112-2112-M-003-020-MY3), and Higher Education Sprout Project of National Taiwan Normal University (NTNU) and the Ministry of Education (MOE) in Taiwan.
@@ Appendix A Direct correspondence between classical and quantum models for $m\leq d$
such that $\sum_{a}\mathcal{I}_{a}$ is trace preserving. We can define a Kraus operator for each non-zero $[T_{a}]_{ij}$, via the action $K_{aij}\left|i\right\rangle=\sqrt{[T_{a}]_{ij}}\left|j\right\rangle$, such that $\mathcal{I}$ acts diagonally in this basis. Now, by choosing $E_{i}=\sigma_{i}=\left|i\right\rangle\left\langle i\right|$, we have a completely dephasing channel $\mathcal{E}_{\text{CD}}$ in the same basis, $\mathcal{E}_{\text{CD}}(\rho):=\sum_{i=1}^{m}\left\langle i\right|\rho\left|i\right\rangle\left|i\right\rangle\left\langle i\right|$. Finally, by using the initial state $\rho_{0}=\sum_{i}\pi_{i}\left|i\right\rangle\left\langle i\right|$, we conclude
$\operatorname{tr}\left[\mathcal{I}_{a_{L}}\circ\mathcal{E}\circ\cdots\circ\mathcal{I}_{a_{1}}\circ\mathcal{E}(\rho_{0})\right]=\pi T_{a_{1}}\cdots T_{a_{L}}\eta.$ (10)
Together with the results from Eqs. 4 and 5, this establishes the correspondence between classical and quantum scenarios for $m\leq d$, and motivates the investigation of the $m>d$ case.
@@ Appendix B Conditions for classicality
Since generic EB channels can violate bounds for classical memories, it is desirable to know the conditions under which non-classical memory effects are actually erased by their application, as often assumed, i.e., such that no choice of quantum instrument $\mathcal{I}$ can lead to non-classical temporal correlations.
A full investigation of these conditions is beyond the scope of this paper. Nevertheless, understanding some sufficient conditions offers insights on the strength of the implicit assumptions being made when applying EB channels as means of characterizing classical memory effects. We present two independent sufficient conditions: either the states $\{\sigma_{i}\}_{i=1}^{m}$ commute, and/or the POVM elements $\{E_{i}\}_{i=1}^{m}$ commute.
If $[\sigma_{i},\sigma_{j}]=0$, $\forall i,j$, then a common eigenbasis exists in which all operators can be written as $\sigma_{i}=\mathrm{diag}(s^{0}_{i},\dots,s^{d-1}_{i})$. Without loss of generality, we can express the channel in this basis as
@@
where $F_{\ell}:=\sum_{i}s_{i}^{\ell}E_{i}$ and $\{F_{\ell}\}_{\ell=1}^{d}$ is a valid POVM. In fact, $s_{i}^{\ell}\geq 0$ since $\sigma_{i}\geq 0$ and $\sum_{\ell}s_{i}^{\ell}=1$ since $\operatorname{tr}\left[\sigma_{i}\right]=1$, which implies $F_{\ell}\geq 0$ and $𝟙\sum_{\ell}F_{\ell}=\sum_{i,\ell}s_{i}^{\ell}E_{i}=\sum_{i}E_{i}=\openone$. We thus have a new definition of the channel $\mathcal{E}$ which uses only $d$ outcomes and states. But from App. A, this implies there exists an equivalent classical model with $d$ states. The case for $[E_{i},E_{j}]=0$ is analogous.
@@ Appendix C Numerical optimization
For $\bm{a}^{L}_{\text{ot}}=0\dots 01$ and $d=L-1$, the best classical bounds known are achieved by the so-called one-way models [6, 25]
$[T_{0}]_{i,i}=1/L,\quad[T_{0}]_{i,i+1}=[T_{1}]_{d,1}=1-1/L,$ (12)
@@
In the quantum case, numerical optimization was performed via gradient descent techniques using the Adam algorithm [40], with an additional exponentially decreasing learning rate from $0.07$ to $10^{-12}$ over 50 000 iterations, and many trials with random initial values. We have found sufficient to optimize for rank-1 $E_{i},\sigma_{i}$, with no further assumptions. Each $\mathcal{I}_{a}$ was defined with a single Kraus operator, as no advantage was observed in using more. As the probability in Eq. 4 is convex for $\rho_{0}$, and we maximize for a single sequence, this initial state can be assumed pure, and w.l.o.g., set to $\rho_{0}=\left|0\right\rangle\left\langle 0\right|$. We have the following constrained optimization problem:
@@
with $\mathcal{E}(\rho)=\sum_{i=1}^{m}\operatorname{tr}\left[\rho E_{i}\right]\sigma_{i}$ and $\mathcal{I}_{a}(\rho)=K_{a}\rho K_{a}^{\dagger}$. This can be relaxed into an unconstrained optimization by using arbitrary $d\times d$ complex matrices $A_{i},B_{i},C_{a}$, then defining $\sigma_{i}:=A_{i}^{\dagger}A_{i}/\mathrm{tr}[A_{i}^{\dagger}A_{i}]$ for the states. For the POVM, we first define $\tilde{B}_{i}:=B_{i}^{\dagger}B_{i}$ and compute $\lambda^{E}_{\text{max}}:=\max\text{eigs}(\sum_{i}\tilde{B}_{i})$, such that letting $E_{i}:=\tilde{B}_{i}/\lambda^{E}_{\text{max}}$ gives $\sum_{i}E_{i}\leq\mathbbm{1}$. For the instrument we compute $\lambda^{\mathcal{I}}_{\text{max}}=\max\text{eigs}(\sum_{a}C_{a}^{\dagger}C_{a})$ and define $K_{a}:=C_{a}/\sqrt{\lambda^{\mathcal{I}}_{\text{max}}}$, giving $\sum_{a}K_{a}^{\dagger}K_{a}\leq\mathbbm{1}$. In this way, the gradient of the objective function will naturally favor solutions satisfying $\sum_{i}E_{i}=\mathbbm{1}$ and $\sum_{a}K_{a}^{\dagger}K_{a}=\mathbbm{1}$. The above construction is general, but can be easily adapted for rank-1 $\sigma_{i}$ and $E_{i}$. The best quantum models found for $L=4$ and $5$ are included in App. D.
Appendix D Numerical examples violating the classical the bound
The following are the best quantum models found for $L=4$ and $5$ using gradient descent techniques, as described in App. C. These optimal values for a given dimension and sequence length required high accuracy numerical optimization to be found, and rely on several significant figures in order to adequately satisfy the constraints in Eq. 13. Despite considerable efforts, no closed-form description of these models was found. Since all optimal $E_{i}$ and $\sigma_{i}$ found were rank-1, we include the vectors such that $E_{i}=\left|e_{i}\right\rangle\left\langle e_{i}\right|$ and $\sigma_{i}=\left|\phi_{i}\right\rangle\left\langle\phi_{i}\right|$. We provide only 4 significant digits of precision, for simplicity, but convergence was obtained up to 10 significant digits. Both cases assume $\rho_{0}=\left|0\right\rangle\left\langle 0\right|$ and $\mathcal{I}_{a}(\rho)=K_{a}\rho K_{a}^{\dagger}$.

### v2

@@ IV Quantum advantages with entanglement-breaking channels
Crucially, the maxima $\Omega_{C}(\bm{a},d)$ must be found explicitly within classical theory, either by exact computation or by obtaining an upper bound, whereas for the quantum value of $p(\bm{a}|\mathcal{I},\mathcal{E},d)$ an explicit model $(\mathcal{I},\mathcal{E})$ must be found. If the classical memory bound has been violated even in the presence of an EB channel $\mathcal{E}$, as in Eq.˜7, then the system still acts as a quantum memory resource for the task of generating temporal correlations.
To achieve this goal, we consider the one-tick sequences [34, 35] of length $L$, defined as $\bm{a}_{\text{ot}}^{L}:=0\dots 01$ and having $\operatorname{DC}(\bm{a}^{L}_{\text{ot}})=L$, under the sub-deterministic scenario d=L1d=L-1. As the case of $L=2$ is trivial, we consider the smallest nontrivial cases $L\geq 3$. For the classical model, analytical upper bounds can be computed via semidefinite programming (SDP) [45]. The case $L=3$ and $4$ were already solved in [45], and here we computed explicitly the case $L=5$ with the same method. In all these cases, the upper bounds computed via SDP coincide, up to the numerical precision, with the values obtained via the one-way models in Refs. [34, 35], which are conjectured to be optimal for all d=L1d=L-1 scenarios. We are therefore sure to have obtained the exact bounds for the cases $L=3,4,$ and $5$. To show EB channels are a quantum memory resource, we must now find an explicit quantum model violating these classical bounds.
One such model can be obtained with the following construction. Consider the deterministic classical model on $m$ states, given by $T_{0}]_{i+1,i}=1$ for $i=1,\dots,m-1$, with $T_{1}]_{1,m}=1$ and zeros for all other entries, and an initial state $\pi=(1,0,\dots,0)^{\text{T}}$. The idea is to approximate this model by emulating $m=d+1$ nearly-orthogonal virtual states with the available $d$-dimensional quantum system. With $\zeta=\mathrm{e}^{2\pi\mathrm{i}/(m-1)}$ and the basis $\{\left|k\right\rangle\}_{k=0}^{d-1}$, let
@@
then define $\mathcal{E}$ with $\sigma_{n}=\left|\psi_{n}\right\rangle\left\langle\psi_{n}\right|$ and $E_{n+1}=\tfrac{d-1}{d}\sigma_{n}$ for $n=1,\dots,m-1$, and finally setting $\rho_{0}=\sigma_{m}=E_{1}=\left|0\right\rangle\left\langle 0\right|$. The instrument is defined as $\mathcal{I}_{a}(\rho)=K_{a}\rho K^{\dagger}_{a}$ with $K_{0}=\mathrm{diag}(1,\cdots,1,0)$ and $K_{1}=\mathbbm{1}-K_{0}$. This quantum model corresponds to isolating the (d1)(d-1)-dimensional subspace orthogonal to $\left|0\right\rangle$, assembling a harmonic complex equiangular tight frame (ETF) [47] of size m1m-1 within this subspace, then using it for the virtual states in the transitions of $T_{0}$. ETFs are useful for this task as they provide a generalization of orthonormal bases, saturating the Welch bounds for minimum overlaps [48]. As seen in Fig.˜3, this quantum ETF construction clearly violates the exact classical bounds for $L=4$ and $5$, conclusively showing that general EB channels do not adequately characterize classical memory effects.
@@
The quantum model presented above provides an explicit violation of the classical bound. This is, however, not necessarily the maximum one can achieve in the quantum case. To investigate the quantum maximum with EB channels, using $m=d+1$, we have also performed numerical optimization via gradient descent methods, obtaining lower bounds for the maxima $\Omega_{Q}(\bm{a},d)=\max_{\mathcal{I},\mathcal{E}}\;p(\bm{a}|\mathcal{I},\mathcal{E},d)$. Results are summarized in Table 1. The quantum models found numerically slightly outperform the previous ETF construction. More details regarding the optimization and the explicit models can be found in appendices Appendices˜C and D, respectively.
V Discussion
Although EB channels are employed as classical memories in a variety of scenarios, the usefulness of a memory ultimately rests in its ability to generate correlations in time. However, a careful accounting of all memory resources is essential to establish a rigorous notion of nonclassicality for temporal correlations, as the dimension of the memory plays a fundamental role [37, 38, 42, 33], and even shared randomness between time steps provides a memory advantage [17, 49, 50, 14]. With this in mind, our results show that it is necessary to look beyond qubits and two-time scenarios when investigating memory effects, as otherwise important distinctions between classical and quantum memories are not apparent [8, 9, 10, 11, 23, 24]. In other words, memory effects should be recognized as fundamentally multi-time phenomena (as also noted in Refs. [14, 51]), and the dimension of all the memory resources must be explicitly taken into account; see also Appendices˜B and D.
The violations of classical bounds, even in the presence of EB channels, reveal that several established notions of classical and quantum memories are still widely misunderstood, arguably due to improper analogies being drawn from spatial correlations. These analogies ignore the particularities of temporal correlations, such as the necessity of bounding classical memory when establishing a notion of nonclassicality, leading to an inadequate treatment of the relevant memory resources. Concretely, while the dimension $m$ of the intermediate classical space of an EB channel plays a fundamental role in its original temporal context—namely, bounding the amount of classical communication between inputs and outputs—, its role is obscured in typical applications of EB channels, where the value of $m$ does not affect the separability of the resulting quantum supermap.
Indeed, in light of our findings, even if unlimited classical communication is considered a free resource, EB channels still cannot be considered “classical memories”. Therefore, the common presumption that EB channels generically act as a classical resource carries additional assumptions, which could be rather strong; see discussion on Appendix˜B. More generally, EB channels should be understood as quantum processes with classical memory [51], while also highlighting their interplay between quantum and classical resources. Our results show this interplay leads to subtle nonclassical memory effects, a phenomenon which warrants further attention.
VI Conclusions
@@
Entanglement-breaking channels are universally recognized as classical resources in a variety of contexts, ranging from quantum networks [8, 9, 10, 23, 11, 24, 52, 25], to multi-time quantum processes and open quantum systems [27, 13, 28, 26, 13, 14, 21]. Quite surprisingly, however, we have found sequential scenarios where they still allow for genuinely quantum memory effects, leading to nonclassical temporal correlations. More precisely, we answer in the negative the following question: is a qudit passing through an entanglement-breaking channel equivalent to a classical dit?
@@
To address this question, we have investigated a task where memory is the central resource, namely, the generation of correlations in time. Indeed, it has been shown that any temporal correlation can be reproduced by a classical system if enough memory is available [41, 42, 43, 44], such that differences between classical and genuinely quantum temporal correlations only emerge if memory is bounded [37, 38, 33, 34, 35]—a fundamental but often neglected feature of temporal correlations. As classical memory is understood in terms of the number of states $d$, a fair comparison between classical and quantum theory must involve a quantum system of the same size. Differences between quantum and classical memories can be found by investigating the most general set of memory operations allowed in each theory, then establishing bounds on their capabilities.
@@
Our findings inspire several future research directions. Understanding the minimal requirements for an EB channel to truly act as a classical memory could offer further insights into the origins of quantum advantages, allowing a more definitive understanding of classical and quantum memory effects; see also Appendix˜A. Furthermore, since the quantum advantage in our case emerges from the repeated interplay between states $\sigma_{i}$ and measurements $E_{i}$, an investigation of the quantum upper bounds in the presence of arbitrary EB channels may inspire new strategies for optimal encoding of classical information within quantum systems in various sequential tasks, with potential applications to, e.g., quantum random-access codes [53, 54, 55, 56, 28], or the quantum compression of classical predictive models for stochastic processes [57, 58, 59]. In addition, even if in this work we have focused on sequential measurements on a single system, our approach for investigating classical and quantum resources via finite-state machines can be naturally extended to arbitrary spatio-temporal scenarios. Indeed, a related result has been recently found in entanglement-assisted communication scenario [60], namely, that the communication of classical messages through a noisy entanglement-breaking qubit channel assisted by quantum entanglement cannot, in general, be simulated by communication through a noisy bit channel assisted by classical correlations. By enabling an exact characterization of memory and communication resources needed for any given task, this approach can lead to novel insights even in simple scenarios, as was the case here.
@@
Naturally, if a quantum memory is required to preserve entanglement, then EB channels are not a useful resource. However, quantum effects and their advantages are not restricted to entanglement [3, 4] or spatial correlations [37, 33, 5, 6, 7]. By recognizing the temporal character inherent to any information processing task, we are required to also consider the fundamental differences between classical and quantum temporal correlations. Our results indicate a more careful and detailed investigation of quantum memory resources is warranted, such that their precise origin and advantage over classical memories can be properly understood, quantified, and successfully applied in emergent quantum technologies.
Acknowledgements
@@
We thank Simon Milz, Philip Taranto, Marco Túlio Quintino and Felix Binder for valuable discussions. This work is supported by the Austrian Science Fund (FWF) through projects ZK 3 (Zukunftskolleg) and F7113 (BeyondC), and partially funded within the QuantERA II Programme that has received funding from the EU’s H2020 research and innovation programme under the GA No 101017733. H.-Y. K. is supported by the National Science and Technology Council (NSTC), (with grant number NSTC 112-2112-M-003- 020-MY3), and the MOE through Higher Education Sprout Project of National Taiwan Normal University.
@@ Appendix A Direct correspondence between classical and quantum models for $m\leq d$
such that $\sum_{a}\mathcal{I}_{a}$ is trace preserving. We can define a Kraus operator for each non-zero $T_{a}]_{ji}$, via the action $K_{aij}\left|i\right\rangle=\sqrt{[T_{a}]_{ji}}\left|j\right\rangle$, such that $\mathcal{I}$ acts diagonally in this basis. Now, by choosing $E_{i}=\sigma_{i}=\left|i\right\rangle\left\langle i\right|$, we have a completely dephasing channel $\mathcal{E}_{\text{CD}}$ in the same basis, $\mathcal{E}_{\text{CD}}(\rho):=\sum_{i=1}^{m}\left\langle i\right|\rho\left|i\right\rangle\left|i\right\rangle\left\langle i\right|$. Finally, by using the initial state $\rho_{0}=\sum_{i}\pi_{i}\left|i\right\rangle\left\langle i\right|$, we conclude
$\operatorname{tr}\left[\mathcal{I}_{a_{L}}\circ\mathcal{E}\circ\cdots\circ\mathcal{I}_{a_{1}}\circ\mathcal{E}(\rho_{0})\right]=\eta T_{a_{L}}\cdots T_{a_{1}}\pi.$ (10)
Together with the results from Eqs.˜4 and 5, this establishes the correspondence between classical and quantum scenarios for $m\leq d$, and motivates the investigation of the $m>d$ case.
@@ Appendix B Conditions for classicality
Since generic EB channels can violate bounds for classical memories, it is desirable to know the conditions under which nonclassical memory effects are actually erased by their application, as often assumed, i.e., such that no choice of quantum instrument $\mathcal{I}$ can lead to nonclassical temporal correlations.
A full investigation of these conditions is beyond the scope of this paper. Nevertheless, understanding some sufficient conditions offers insights on the strength of the implicit assumptions being made when applying EB channels as means of characterizing classical memory effects. Assuming input and output spaces have the same number of orthogonal states, we present two independent sufficient conditions: either the states $\{\sigma_{i}\}_{i=1}^{m}$ commute, and/or the POVM elements $\{E_{i}\}_{i=1}^{m}$ commute.
If $\sigma_{i},\sigma_{j}]=0$, $\forall i,j$, then a common eigenbasis exists in which all operators can be written as $\sigma_{i}=\mathrm{diag}(s^{0}_{i},\dots,s^{d-1}_{i})$. Without loss of generality, we can express the channel in this basis as
@@
[…]
where $\tilde{\sigma}_{\ell}:=\sum_{i=1}^{m}e_{i}^{\ell}\sigma_{i}$ are valid states, since $e_{i}^{\ell}\geq 0$ and $\sum_{i=1}^{m}e_{i}^{\ell}=1$ for every $\ell$, and we once again obtain a channel $\mathcal{E}$ written in terms of $d$ outcomes.
The above results hold under the assumption that input and output spaces of $\mathcal{E}$ have the same dimension. To understand why this is necessary, consider two separate EB channels, the quantum-classical $\mathcal{E}_{1}(\rho)=\sum_{i=1}^{m}\operatorname{tr}\left[\rho E_{i}\right]\left|i\right\rangle\left\langle i\right|$ and the classical-quantum $\mathcal{E}_{2}(\rho)=\sum_{i=1}^{m}\operatorname{tr}\left[\left|i\right\rangle\left\langle i\right|\rho\right]\sigma_{i}$, such that $\mathcal{E}=\mathcal{E}_{2}\circ\mathcal{E}_{1}$. The output states of $\mathcal{E}_{1}$ and POVM elements of $\mathcal{E}_{2}$ are both $\{\left|i\right\rangle\left\langle i\right|\}_{i=1}^{m}$, which are completely classical and within an $m$-dimensional space. Thus, taken individually, it would seem that both $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$ should be completely classical. Nevertheless, their concatenation leads to nonclassical temporal correlations, as we have shown. The fact $m>d$ has allowed subtle quantum effects to emerge from the interplay between sequential encoding and decoding of quantum information into the classical memory. The sequential nature of our scenario prevents a straightforward analysis, however, as nonclassical effects only emerge beyond qubits and at least four time steps.
We also note that these conditions for classicality are reminiscent to those for zero quantum discord in bipartite separable states [61], suggesting that non-zero discord of the Choi state of an EB channel may play a role in its temporal nonclassicality. We leave such investigations to future research.
@@ Appendix C Numerical optimization
For $\bm{a}^{L}_{\text{ot}}=0\dots 01$ and d=L1d=L-1, the best classical bounds known are achieved by the so-called one-way models [33, 35]
$T_{0}]_{i,i}=1/L,\quad[T_{0}]_{i,i+1}=[T_{1}]_{d,1}=1-1/L,$ (13)
@@
In the quantum case, numerical optimization was performed via gradient descent techniques using the Adam algorithm [62], with an additional exponentially decreasing learning rate from $0.07$ to $10^{-12}$ over 50 000 iterations, and many trials with random initial values. We have found sufficient to optimize for rank-1 $E_{i}$ and $\sigma_{i}$, with no further assumptions. Each $\mathcal{I}_{a}$ was defined with a single Kraus operator, as no advantage was observed in using more. As the probability in Eq.˜4 is convex for $\rho_{0}$, and we maximize for a single sequence, this initial state can be assumed pure, and w.l.o.g., set to $\rho_{0}=\left|0\right\rangle\left\langle 0\right|$. We have the following constrained optimization problem:
@@
with $\mathcal{E}(\rho)=\sum_{i=1}^{m}\operatorname{tr}\left[\rho E_{i}\right]\sigma_{i}$ and $\mathcal{I}_{a}(\rho)=K_{a}\rho K_{a}^{\dagger}$. This can be relaxed into an unconstrained optimization by using arbitrary $d\times d$ complex matrices $A_{i},B_{i},C_{a}$, then defining $\sigma_{i}:=A_{i}^{\dagger}A_{i}/\mathrm{tr}[A_{i}^{\dagger}A_{i}]$ for the states. For the POVM, we first define $\tilde{B}_{i}:=B_{i}^{\dagger}B_{i}$ and compute $\lambda^{E}_{\text{max}}:=\max\text{eigs}(\sum_{i}\tilde{B}_{i})$, such that letting $E_{i}:=\tilde{B}_{i}/\lambda^{E}_{\text{max}}$ gives $\sum_{i}E_{i}\leq\mathbbm{1}$. For the instrument we compute $\lambda^{\mathcal{I}}_{\text{max}}=\max\text{eigs}(\sum_{a}C_{a}^{\dagger}C_{a})$ and define $K_{a}:=C_{a}/\sqrt{\lambda^{\mathcal{I}}_{\text{max}}}$, giving $\sum_{a}K_{a}^{\dagger}K_{a}\leq\mathbbm{1}$. In this way, the gradient of the objective function will naturally favor solutions satisfying $\sum_{i}E_{i}=\mathbbm{1}$ and $\sum_{a}K_{a}^{\dagger}K_{a}=\mathbbm{1}$. The above construction is general, but can be easily adapted for rank-1 $\sigma_{i}$ and $E_{i}$. The best quantum models found for $L=4$ and $5$ are included in Appendix˜D, below.
Appendix D Numerical examples violating the classical bound
The following are the best quantum models found for $L=4$ and $5$ using gradient descent techniques, as described in Appendix˜C. These optimal values for a given dimension and sequence length required high accuracy numerical optimization to be found, and rely on several significant figures in order to adequately satisfy the constraints in Eq.˜14. Despite considerable efforts, no closed-form description of these models was found. Since all optimal $E_{i}$ and $\sigma_{i}$ found were rank-1, we include the vectors such that $E_{i}=\left|e_{i}\right\rangle\left\langle e_{i}\right|$ and $\sigma_{i}=\left|\phi_{i}\right\rangle\left\langle\phi_{i}\right|$. We provide only 5 significant digits of precision, for simplicity, but convergence was obtained up to 10 significant digits. Both cases assume $\rho_{0}=\left|0\right\rangle\left\langle 0\right|$ and $\mathcal{I}_{a}(\rho)=K_{a}\rho K_{a}^{\dagger}$. For $L=3$ and $d=2$, we were unable to find any violation of the classical bound, strongly suggesting EB-based temporal nonclassicality only emerges beyond two-level systems. Furthermore, since $\operatorname{DC}(\bm{a})\leq L$ for any sequence of length $L$, either a qubit or a bit are sufficient to produce any correlations on two time steps. It is therefore essential to also look beyond two-time steps in order to reveal fundamental differences between classical and quantum systems in this sequential measurement scenario.